# Bridging Cognitive Maps: a Hierarchical Active Inference Model of Spatial Alternation Tasks and the Hippocampal-Prefrontal Circuit

### v2

@@ 4.1 Clone-structured cognitive graphs (CSCG) for learning cognitive maps of physical and task space
Figure 6: Learned cognitive graphs. The arrows indicate possible transitions from one state to the next. The node in the center where all nodes point to, is the added and absorbing “dummy” state.
(a) The learned CSCG for the physical space model (level 1). The color in the graph represents distinct observations, only the states that are active when pursuing the spatial alternation rule (LCRC) are colored, the other states are marked in gray. (b) The learned CSCG for the task space model (level 2). The color in the graph indicates the states that were active when pursuing the spatial alternation rule (LCLR). While only four states are necessary to encode the rule, it can be observed that many more states are added by learning the other paths that might have been visited during training.
@@ 4.2.2 Active inference
$G(\pi,\tau)\geq-\underbrace{\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}\big{[}D_{KL}[Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi)||Q(\mathbf{s}_{\tau}|\pi)]\big{]}}_{\text{Epistemic value}}-\underbrace{\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}\big{[}\log P(\mathbf{o})\big{]}}_{\text{Pragmatic Value}}$ (4)
This quantity is decomposed into an epistemic value and a pragmatic value (Friston et al., 2017a, ). The epistemic value computes the expected information gain (Bayesian surprise) between the prior $Q(\mathbf{s}_{\tau}|\pi)$ and posterior $Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi)$. Intuitively, this quantity represents how much the agent expects the belief over the state to shift when pursuing this policy. The pragmatic value then measures the expected log probability of observing the preferred observation under the selected policy, intuitively computing how likely it is that this policy will drive the agent to its prior preferences. The full expected free energy for a finite time horizon $T$ is computed as $\sum_{\tau=1}^{T}G(\pi,\tau)$.
@@ 4.3 Casting CSCGs as partially observable Markov decision processes
Figure 8: Factor graph of a POMDP: The conditional dependencies in Hidden Markov Models are parameterized by a set of matrices. The $\mathbf{A}$ matrix parameterizes the likelihood model, i.e. how states $\mathbf{s}_{t}$ map to observations $\mathbf{o}_{t}$. The $\mathbf{B}$ matrix parameterizes the transition model, i.e. how state $\mathbf{s}_{t+1}$ changes at each timestep, dependent on policy $\mathbf{\pi}$. The $\mathbf{C}$ vector denotes the preferred observation or state, and directly influences the $\mathbf{G}$ factor which conditions $\pi$. The policy $\pi$ depends on the habit $E$
. Orange circles denote that the variable is observed.
@@ Acknowledgements
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Specific Grant Agreements No. 945539 (Human Brain Project SGA3) and No. 952215 (TAILOR); the European Research Council under the Grant Agreement No. 820213 (ThinkAhead), the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, “FAIR”; Project PE0000006, “MNESYS”), and the PRIN PNRR P20224FESY. The GEFORCE Quadro RTX6000 and Titan GPU cards used for this research were donated by the NVIDIA Corporation. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@ Appendix A Learning the spatial level using a Hidden Markov Model
We investigate the benefit of using the CSCG’s for learning the distinct layers of the hierarchy over using standard Hidden Markov Models on a navigation task. Specifically, we learn the spatial level using both approaches on the same dataset: a trajectory of the agent that contains a path from each distinct pose (position and orientation) to another. We consider the following variants, where the model is used by an active inference agent implemented in pymdp Heins et al., (2022): (i) a CSCG agent, as used in the main text, (ii) a HMM agent trained using the Baum-Welch algorithm (on the same data), and (iii) a HMM fine tuned by the Viterbi algorithm (on the same data).

### v3

@@ 4.1 Clone-structured cognitive graphs (CSCG) for learning cognitive maps of physical and task space
Figure 6: Learned cognitive graphs. The arrows indicate possible transitions from one state to the next. The node in the center where all nodes point to, is the added and absorbing “dummy” state. (a) The learned CSCG for the physical space model (level 1). The color in the graph represents distinct observations, only the states that are active when pursuing the spatial alternation rule (LCRC) are colored, the other states are marked in gray. (b) The learned CSCG for the task space model (level 2). The color in the graph indicates the states that were active when pursuing the spatial alternation rule (LCLR). While only four states are necessary to encode the rule, it can be observed that many more states are added by learning the other paths that might have been visited during training.
@@
The CSCGs are learned using the expectation-maximization (EM) algorithm for HMMs (the Baum-Welch algorithm) which maximizes the ELBO as described in George et al., (2021). During the E-step, the posteriors over states are estimated through smoothing, i.e. the forward-backward algorithm. The M-step then selects the optimal parameters for the transition model, given this sequence of visited states. For the update equations, the reader is referred to George et al., (2021).
@@ 4.2.2 Active inference
$G(\pi,\tau)$ $=-\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}\big{[}D_{KL}[Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi)||Q(\mathbf{s}_{\tau}|\pi)]\big{]}-\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}\big{[}\log P(\mathbf{o})\big{]}$ (4)
$\quad+\underbrace{\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}[D_{KL}(Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau})||P(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi))]}_{\text{Expected Approximation Error}\geq 0}$ (5)
$\geq-\underbrace{\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}\big{[}D_{KL}[Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi)||Q(\mathbf{s}_{\tau}|\pi)]\big{]}}_{\text{Epistemic value}}-\underbrace{\mathbb{E}_{Q(\mathbf{o}_{\tau}|\pi)}\big{[}\log P(\mathbf{o})\big{]}}_{\text{Pragmatic Value}}$ (6)
This quantity is decomposed into an epistemic value and a pragmatic value (Friston et al., 2017a, ). When evaluating $G$, we compute the upper bound shown in Equation (6). We thus make the assumption that the expected approximation error (Equation (5)) is zero (Heins et al.,, 2022). The epistemic value computes the expected information gain (Bayesian surprise) between the prior $Q(\mathbf{s}_{\tau}|\pi)$ and posterior $Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi)$. Intuitively, this quantity represents how much the agent expects the belief over the state to shift when pursuing this policy. The pragmatic value then measures the expected log probability of observing the preferred observation under the selected policy, intuitively computing how likely it is that this policy will drive the agent to its prior preferences. The full expected free energy for a finite time horizon $T$ is computed as $\sum_{\tau=1}^{T}G(\pi,\tau)$.
@@ 4.3 Casting CSCGs as partially observable Markov decision processes
Figure 8: Factor graph of a POMDP: The conditional dependencies in Hidden Markov Models are parameterized by a set of matrices. The $\mathbf{A}$ matrix parameterizes the likelihood model, i.e. how states $\mathbf{s}_{t}$ map to observations $\mathbf{o}_{t}$. The $\mathbf{B}$ matrix parameterizes the transition model, i.e. how state $\mathbf{s}_{t+1}$ changes at each timestep, dependent on policy $\mathbf{\pi}$. The $\mathbf{C}$ vector denotes the preferred observation or state, and directly influences the $\mathbf{G}$ factor which conditions $\pi$. The policy $\pi$ depends on the habit $E$. Orange circles denote that the variable is observed.
@@
Using these tensors, inference at each single level can be implemented as a Bayes filter, iteratively computing the posterior over state at each timestep using the following update:
$\mathbf{qs}_{t}=\sigma(\mathbf{A}\cdot\mathbf{o}_{t}+\mathbf{B}_{\pi_{t-1}}\cdot\mathbf{qs}_{t-1}),$ (9)
where $\mathbf{qs}_{t}$ denotes the parameters of the categorical distribution over state, i.e. $Q(\mathbf{s}_{\tau}|\mathbf{o}_{\tau},\pi)=\text{Cat}(\mathbf{qs}^{1}_{\tau})$, $\mathbf{o}_{t}$ represents the observation as a one-hot vector, $\sigma$ denotes the softmax function, and the $\mathbf{B}$-tensor is sliced by the policy $\pi_{t-1}$ (Friston et al., 2017a, ). In the first timestep, $\mathbf{qs}_{0}$ is initialized as the prior matrix $\mathbf{D}$.
The expected free energy $G$ for a considered policy $\pi_{t}$ can then be evaluated using these tensors for each future timestep $\tau$. As we specify the preference over state for the first level, and over observation for the second level, the distinction is made explicitly. For the first level, this boils down to (Parr et al.,, 2022):
$G^{1}_{\tau}$ $=\text{diag}(\mathbf{A}^{1}{}^{\top}\cdot\ln\mathbf{A}^{1})\cdot\mathbf{qs}^{1}_{\tau}-\mathbf{qo}^{1}_{\tau}\cdot\ln\mathbf{qo}^{1}_{\tau}+\mathbf{qs}^{1}_{\tau}\cdot\ln\mathbf{C}^{1},$ (10)
where the superscript 1 denotes the first level, $\mathbf{qs}^{1}_{\tau}$ denotes the parameters of the categorical distribution over state, i.e. $Q(\mathbf{s}^{1}_{\tau}|\mathbf{o}^{1}_{\tau},\pi^{1}_{t})=\text{Cat}(\mathbf{qs}^{1}_{\tau})$, which is computed as $\mathbf{qs}^{1}_{\tau}=\mathbf{B}_{1,\pi^{2}_{t}}\cdot\mathbf{qs}^{1}_{t}$. The parameters of the categorical over the observation is denoted by $\mathbf{qo}^{1}_{\tau}$, i.e. $Q(\mathbf{o}^{1}_{\tau})=\text{Cat}(\mathbf{qo}^{1}_{\tau})$, and is computed as $\mathbf{qo}^{1}_{\tau}=\mathbf{A}^{1}\cdot\mathbf{qs}^{1}_{\tau}$. The same equation can be used for the expected free energy at the second level, except now the preference is specified as an expected observation:
$G^{2}_{\tau}$ $=\text{diag}(\mathbf{A}^{2}{}^{\top}\cdot\ln\mathbf{A}^{2})\cdot\mathbf{qs}^{2}_{\tau}-\mathbf{qo}^{2}_{\tau}\cdot\ln\mathbf{qo}^{2}_{\tau}+\mathbf{qo}^{2}_{\tau}\cdot\ln\mathbf{C}^{2},$ (11)
where the superscript 2 denotes the second level. $Q(\mathbf{s}^{2}_{\tau}|\mathbf{o}^{2}_{\tau},\pi^{1}_{t})=\text{Cat}(\mathbf{qs}^{2}_{\tau})$, which is computed as $\mathbf{qs}^{2}_{\tau}=\mathbf{B}_{2,\pi^{2}_{t}}\cdot\mathbf{qs}^{2}_{t}$. The parameters of the categorical over the observation at level 2 is denoted by $\mathbf{qo}^{2}_{\tau}$, i.e. $Q(\mathbf{o}^{2}_{\tau})=\text{Cat}(\mathbf{qo}^{2}_{\tau})$, and is computed as $\mathbf{qo}^{2}_{\tau}=\mathbf{A}^{2}\cdot\mathbf{qs}^{2}_{\tau}$. The full expected free energy $G$ can then be acquired by summing over time horizon $T$: $G=\sum_{\tau=t+1}^{T}G_{\tau}$, and the posterior over the policies is achieved through $Q(\pi)=\sigma(\gamma\cdot G+\ln E)$.
@@ Acknowledgements
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Specific Grant Agreements No. 945539 (Human Brain Project SGA3) and No. 952215 (TAILOR); the European Research Council under the Grant Agreement No. 820213 (ThinkAhead), the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, CUP B53C22003630006, ”FAIR”; Project PE0000006, CUP J33C22002970002 “MNESYS”), and the PRIN PNRR P20224FESY. The GEFORCE Quadro RTX6000 and Titan GPU cards used for this research were donated by the NVIDIA Corporation. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@ Appendix A Learning the spatial level using a Hidden Markov Model
We investigate the benefit of using the CSCG’s for learning the distinct layers of the hierarchy over using standard Hidden Markov Models on a navigation task. Specifically, we learn the spatial level using both approaches on the same dataset: a trajectory of the agent that contains a path from each distinct pose (position and orientation) to another. We consider the following variants, where the model is used by an active inference agent implemented in pymdp (Heins et al.,, 2022): (i) a CSCG agent, as used in the main text, (ii) a HMM agent trained using the Baum-Welch algorithm (on the same data), and (iii) a HMM fine tuned by the Viterbi algorithm (on the same data).